OpenStudy (anonymous):

I need some help with a few word problems. -A cannery processed 1,445 pounds of strawberries in 4.5 hours. The cannery processed 2,320 pounds in 7 hours. Write a linear equation to model the relationship between the weight of the strawberries S and time T. How many pounds of strawberries can be processed in 11 hours? Thank you in advance for your help!

hero (hero):

Hint. (4.5, 1445) and (7, 2320) are points. Find the slope. Then find the y-intercept , b. Then use y = mx + b to write the equation of the line.

OpenStudy (anonymous):

so m=875/2.5? in y=mx+b y=875/2.5x+11?

hero (hero):

Show the steps you used to get that please.

OpenStudy (anonymous):

\[m=\frac{ 2320-1445 }{ 7-4.5 }\] \[m=\frac{ 875 }{ 2.5}\] \[y=\frac{ 875 }{ 2.5 }x+11\] 11 being the number of hours. right?

OpenStudy (anonymous):

Are those the correct steps?

hero (hero):

You didn't solve for b properly. \(x\) represents hours \(y\) represents pounds of strawberries. To solve for b, you need to input the slope and a point in to the formula: y = mx + b Assuming your slope is correct, you input one of the points \((x,y) = \color\green{(4.5, 1445)}\) and the slope \(m = \color\red{(875/2.5)}\) \(\color\green{1445} = \color\red{(875/2.5)}\color\green{(4.5)} + b\)

hero (hero):

Then solve for \(b\). Once you have solved for \(b\). Then you can re-write your equation of the line but only include the values of \(m\) and \(b\)